OpenStudy (anonymous):

find the equation of the line through the points(1,-3) and parallel to the lines through (1,2) and (4,3)

OpenStudy (anonymous):

line that goes through (1,2) and (4,3): y=mx+a m=1/3 Since it is parallel, the slope should be same. \[(y+3)=\frac{1}{3}(x-1)\] \[y=\frac{x}{3}-\frac{10}{3}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome!